A lock on a bank vault has three dials each with 30 positions. For the vault to open, each dial must be in the correct position. What is the probability that if you randomly select a combination, you will be able to open the vault (report this as the whole number denominator of 1/x)?
Answer #1

total number of combinations on 3 dials =30*30*30 =27000

probability that if you randomly select a combination, you will be able to open the vault =1/27000